OpenStudy (goldps):

how do i find the zeros of f(x)=3x^3+2

OpenStudy (amistre64):

let f(x) = 0

OpenStudy (amistre64):

the rest is pretty basic algebra

OpenStudy (amistre64):

you want all 3 roots, real and complex?

OpenStudy (goldps):

ok let me try that

OpenStudy (goldps):

i got 2 and that wasnt right, the zeros need to cross the x axis

OpenStudy (amistre64):

f(x)=3x^3+2 0 = 3x^3 + 2 -2 = 3x^3 -2/3 = x^3 cbrt(-2/3) = x

OpenStudy (goldps):

ok got it